OpenStudy (anonymous):

Convert to polar form xy= 6

OpenStudy (turingtest):

conversion from cartesian to polar:\[x=r\cos\theta\]\[y=r\sin\theta\]

OpenStudy (anonymous):

then what

OpenStudy (turingtest):

then substitute\[r\cos\theta\cdot r\sin\theta=6\]and simplify you may want to use some trig identities

OpenStudy (turingtest):

@Busterkitten look, it's a simple substitution^^^

OpenStudy (anonymous):

what did you get for the substitution

OpenStudy (asnaseer):

Any point P can be represented in cartesian coordinates as follows:  In this system, the coordinates of P are (x,y)

OpenStudy (asnaseer):

In Polar coordinates, P can also be represented as its distance (r) from the origin and the angle (\(\theta\)) that it makes with the x-axis: